The commercial division of a real estate firm is conducting a regression analysis of the relationship between annual gross reIs this right?, if not, please answer A-E correctly. Thanks
Answer #1

a) number of apartment buildings =9

b) yhat =20.00+7.21x

c)p value =0.0011

Conclusion: there is a significant linear relationship

d)F =41587.2/((51984.3-41587.2)/7)=28.00

p value =0.0011

Conclusion: there is a significant linear relationship

e)

predicted value =20+7.21*50 =380.5